What is the measurement of the supplement if the measure of the angle is eleven times greater than its supplement?

Answer by jim_thompson5910(35256)

11x + x = 180

12x = 180

x = 180/12

x = 15

11x = 11*15

11x = 165

So the angle is 165 degrees and the supplement is 15 degrees.